OpenStudy (anonymous):

Write the function in vertex form, and identify its vertex. g(x) = 5x2 - 50x + 128

OpenStudy (anonymous):

first coordinate of the vertex is always \[-\frac{b}{2a}\] in your case you get \[-\frac{-50}{2\times 5}=10\] so that is the first coordinate

OpenStudy (anonymous):

well that was wrong wasn't it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{-50}{2\times 10}=5\] my mistake

OpenStudy (anonymous):

so first coordinate of the vertex is 5, to find the second \g(x) = 5x2 - 50x + 128\] \[g(5)=5\times 5^2-50\times 5+128=3\] so vertex is (5,3)

OpenStudy (anonymous):

then to write in "vertex" form put \[g(x)=5(x-5)^2+3\]